OpenStudy (anonymous):

Chameleons catch insects with their tongues, which they can rapidly extend to great lengths. In a typical strike, the chameleon's tongue accelerates at a remarkable 280 for 20 , then travels at constant speed for another 30 .

OpenStudy (anonymous):

280 m/s for 20 s?

OpenStudy (anonymous):

280 m/s2 for 20 ms During this total time of 50ms , 1/20 of a second, how far does the tongue reach?

OpenStudy (anonymous):

so you would have \[(280-0)/20 = 14 \] for the average initial speed. Then\[280*30=8400\] so 8400 + 14 *20

OpenStudy (anonymous):

First we find v_1 which is a*t_1 a= 280m/s^2 t_1=0.02s --------- Then we know it travels with an average speed of v_1/2=v_2 during it's acceleration period. --------- We know it travels at v_1 for a period of t_2 t_2 = 0.03 s --------- Total distance: s = v_2*t_1+v_1*t_2

OpenStudy (anonymous):

Leaving us with the answer s = 0.224 m = 22.4 cm